OpenStudy (anonymous):

whats sine and cosine of 5pie/4 in unit circle? whats sine and cosine of 5pie/4 in unit circle? @Mathematics

OpenStudy (anonymous):

mmm pie

OpenStudy (anonymous):

trigonometric function on unit circle

OpenStudy (anonymous):

find \[\frac{5\pi}{4}\] on the unit circle on the last page of cheat sheet. first coordinate is \[\cos(\frac{5\pi}{4})\] second one is \[\sin(\frac{5\pi}{4})\]

OpenStudy (anonymous):

btw, it is just the trig function of a number, not "on the unit circle" although that is how you would find them. btw btw it is "pi" greek letter not "pie" as in cherry

OpenStudy (anonymous):

lol help me out,i dont know how to determine them on the unit circle

OpenStudy (asnaseer):

 the diagram may help explain it...